OpenStudy (anonymous):

integral of (x^6ln(x))dx

OpenStudy (anonymous):

Integrate by parts. Like 5 or 6 times.

OpenStudy (anonymous):

derive lnx and integrate x^6, referring to integration by parts.

OpenStudy (anonymous):

\[\Large \int x^6\ln x dx = \frac{1}{7}x^7\ln x - \frac{1}{7}\int x^7 \cdot \frac{1}{x} \]

OpenStudy (anonymous):

I take back what I said! Overthinking, yet again...

OpenStudy (anonymous):

@SithsAndGiggles , you wouldn't believe how often I used integration by parts when a regular substitution would have saved me 30 minutes of work (x !

OpenStudy (anonymous):

Tch, happens to the best of us.

OpenStudy (anonymous):

I figured it out :) thanks guys!